OpenStudy (desmarie):

Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.)

OpenStudy (desmarie):

the answer is 9/2 pi - 9 but how?

OpenStudy (ineedhelplz):

diameter of the circle is the diagonal of the square. since the square has side 6, we know the diameter of the circle is \[d=6\sqrt{2}\] and so the radius is \[3\sqrt{2}\] area of square is 36 area of circle is \[\pi r^2=\pi (3\sqrt{2})^2=18\pi\]

OpenStudy (ineedhelplz):

so we know the area of the square, and we know the area of the circle. if we knew what 'ONE SEGMENT' meant i we would be done

OpenStudy (ineedhelplz):

\[\frac{18(\pi-2)}{4}=\frac{9(\pi-2)}{2}\]